OpenStudy (anonymous):

will fan and medal Find the curvature of r(t) = (12sint,5t,12cost) at the point (0,5pi,-12)

OpenStudy (anonymous):

i used second formula in http://tutorial.math.lamar.edu/Classes/CalcIII/Curvature_files/eq0005P.gif

OpenStudy (astrophysics):

Yeah, that should work

OpenStudy (anonymous):

i have an answer and would like someone to verify my answer is correct

OpenStudy (astrophysics):

Sure, can you please show all your steps though.

OpenStudy (anonymous):

my answer came out to 12

OpenStudy (astrophysics):

Well that's not very useful to me, if you want to check your steps post them.

OpenStudy (anonymous):

i dont want to type it all out. i was hoping someone would work the problem to check

OpenStudy (astrophysics):

We're not here to do your work, but we are willing check YOUR work.

OpenStudy (anonymous):

ok ill wait for someone who is capable of doing the problem

OpenStudy (anonymous):

What makes you think anyone here is willing to type it all out? At the very least, you could show the tiniest bit of effort to perhaps *show* you've tried.

OpenStudy (anonymous):

i do not anyone to type it out. Just to work the problem. i am not a very fast typer and suck with the equation button. i showed the formula i was using

OpenStudy (anonymous):

its an exam review with no answers. I was hoping someone would work the problem and say yes that is what i got or no that is not what i got

zepdrix (zepdrix):

Grr I'm a lil rusty on these. What is T again? \[\Large\rm \hat T=\frac{\vec r'}{||\vec r'||}\]Like that or something?

OpenStudy (anonymous):

Tangent

zepdrix (zepdrix):

Mmm ok I'll try using that one then. >.< In case we end up with different answers, then we can maybe back up and see where we went wrong.

OpenStudy (anonymous):

ok sounds great

zepdrix (zepdrix):

Might take me a few minutes >.< ahhh the rust is real

OpenStudy (anonymous):

Checking with Mathematica, the answer is not \(12\).

OpenStudy (astrophysics):

If you can show your steps, you know we can pinpoint where you made the mistake...otherwise seems you just want a direct answer.

OpenStudy (anonymous):

that strange that i would get perfect squares in my simplification then. What was it?

OpenStudy (anonymous):

ohh wait i got my mistake

zepdrix (zepdrix):

\[\Large\rm \vec r=\left<12 \sin t, 5t, 12 \cos t\right>\]Then,\[\Large\rm \vec r'=\left<12 \cos t, 5, -12 \sin t\right>\]and,\[\Large\rm \vec r''=\left<-12 \sin t, 0, -12\cos t\right>\] And you were using the second method?

OpenStudy (anonymous):

i did not cube the bottom was it 12/169 ?

OpenStudy (anonymous):

yes i was using the second method and that is what i got for the r values

OpenStudy (anonymous):

@SithsAndGiggles was it 12/169

OpenStudy (anonymous):

No, but it's certainly closer. You're off by about 0.03. Have you attempted using the formula on the left?

OpenStudy (anonymous):

if im off by .03 then maybe mathmatica is a approximation instead of exact? i am not using a calculator

zepdrix (zepdrix):

Yah I hate doing that stupid cross product >.<

OpenStudy (anonymous):

i agree lol

OpenStudy (anonymous):

\[\begin{align*}\vec{T}(t)&=\frac{\vec{r}~'(t)}{\|\vec{r}~'(t)\|}\\\\&=\frac{\langle 12\cos t,5,-12\sin t\rangle}{\sqrt{12^2\cos^2t+25+(-12)^2\sin^2t}}\\\\&=\frac{\langle 12\cos t,5,-12\sin t\rangle}{\sqrt{144+25}}\\\\ &=\frac{1}{13}\langle 12\cos t,5,-12\sin t\rangle\end{align*}\] Is this what you have?

OpenStudy (anonymous):

i never calculated the tangent but yeah that looks right

zepdrix (zepdrix):

Hmm I came up with 12/169 also :P darn

OpenStudy (anonymous):

i think that is probably right then. did you do the first formula?

zepdrix (zepdrix):

Ya, I kept screwing up doing that cross product, so I switched to the first one. Wolfram seems to agree. :O https://www.wolframalpha.com/input/?i=curvature+of+r%3D%3C12+sin+t%2C+5t%2C+12+cos+t%3E+at+t%3Dpi

OpenStudy (anonymous):

wow thanks i did know how to type that in to wolf, that would of saved a lot of time

OpenStudy (anonymous):

thanks for the help

zepdrix (zepdrix):

Ya, whenever you're not quite sure how to type it, try using a lot of "words" instead of symbols. usually works :O

OpenStudy (anonymous):

ok i tried typing the vector into it but i never typed curvature. good to know. thank as well sith

zepdrix (zepdrix):

Aw Siff >.< you were gonna show up the magic lol

zepdrix (zepdrix):

Doing both methods, I didn't see a time where I could plug in the t value. Every t just fell out in the end. So this surface has constant curvature I guess? It's some type of sphericalllll shinanigans or something?

OpenStudy (astrophysics):

Yes, we often call it an osculating sphere if k>0

OpenStudy (anonymous):

does that look right to you astro lol

OpenStudy (anonymous):

There might be some typo in the Mathematica file then. I'm getting \(\dfrac{12}{169}\) now. We have \(\vec{T}~'(t)=\dfrac{1}{13}\langle -12\sin t,0,-12\cos t\rangle\), which means \(\left\|\vec{T}~'(t)\right\|=\sqrt{\dfrac{(-12)^2\sin^2t+0^2+(-12)^2\cos^2t}{13^2}}=\dfrac{12}{13}\). Then\[\begin{align*} \kappa(t)&=\frac{\dfrac{12}{13}}{13}=\frac{12}{169} \end{align*}\]

OpenStudy (astrophysics):

Yes it's good

OpenStudy (anonymous):

good i glad your here to say you agree

OpenStudy (astrophysics):

Sorry to doubt you I suppose, I deal with a lot of people who just want the answer.

OpenStudy (anonymous):

@zepdrix Yeah the curve defines a helix.

zepdrix (zepdrix):

Hmm neato :)

zepdrix (zepdrix):

@SithsAndGiggles @Astrophysics I dunno why you guys fuss about it so much :D If someone doesn't care enough about their education to actually try, it's on them. It'll catch up to them eventually. They're just doing themselves a disservice. I guess I just like to assume that people have good intentions when they're on this site, maybe that's naive though :)

OpenStudy (astrophysics):

Thanks @zepdrix I often tell myself that to, you have the right idea, they will eventually learn I guess, the way I handle it never really works or it's very rare. Thanks :)

OpenStudy (anonymous):

this is calculus 3 as well, if i just put the answer down my prof would laugh and give me a zero even if it was correct